OpenStudy (anonymous):

an advertisement claims that a 5-pound box of detergent will wash 20 loads of laundry a 5 pound box cost $3.50 a 15 pound box cost $10.00 how much will you save per load if you buy a 15 pound box

OpenStudy (campbell_st):

5 pound box: cost per wash = 3.50/20 = $0.175 per wash 15 pound box: cost per wash = 10/(3 x 20) = $0.167 per wash the 15 pound wash will save 0.175 - 0.167 = $0.008 per wash